OpenStudy (anonymous):

How many different atoms are there in HNO3? 3 or 5 or what?

OpenStudy (anonymous):

3, because there are 3 elements (there are 5 atoms, but only 3 DIFFERENT kinds of atoms)

OpenStudy (anonymous):

So what about Cu(NO3)2?

OpenStudy (anonymous):

try it: how many different elements are there? (that's also the number of different atoms)

OpenStudy (anonymous):

There's three diff. atoms

OpenStudy (anonymous):

correct! :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So what about the numbers? Should we do anything with them?

OpenStudy (anonymous):

no, when you are triing to find the number of different atoms :) each element has a unique atom and that's it :)

OpenStudy (anonymous):

Cool, I appreciate the help :D

OpenStudy (anonymous):

Would you mind if you could help with one more example?

OpenStudy (anonymous):

C12 H22 O11, that's just 3 atoms?

OpenStudy (anonymous):

3 DIFFERENT ones :) (but the total is 12+22+11=45)

OpenStudy (anonymous):

Cool :)

OpenStudy (anonymous):

whenever you need help, just tag me :)

OpenStudy (anonymous):

Are you able to balance equations?

OpenStudy (anonymous):

Chemical equations?

OpenStudy (anonymous):

Yeah sure I'll tag you :)

OpenStudy (anonymous):

this might help you :) http://www.wikihow.com/Balance-Chemical-Equations

OpenStudy (anonymous):

Gee you're really helpful

OpenStudy (anonymous):

Btw, do you know the difference between the modes on a scientific calculator?

OpenStudy (anonymous):

SD - standard deviation calculations COMP - normal calculating DEG - calculating using degrees RAD - calculating using radians GRA - calculating using grads FIX - number of decimal place specifications SCI - number of significant digit specification NORM - cancels FIX and SCI settings I had to look to the calculator manual, but I can explain all of them, if you want :)

OpenStudy (anonymous):

Just DEG, RAD and GRA

OpenStudy (anonymous):

you probably know, what is a degree. 1 radian is 2*pi degrees. (1 rad =3,14*2=6,28° ) 1 gradian is radian/400. So: 1 rad = 400 grad 1 grad = 1/400 = 0,0025 rad 1rad = 6,28 deg 1 deg = 1/6,28 rad 1 grad = 1/6,28/400 rad 1 deg = 1*6,28*400 grad